How many pounds of a 15% aluminum alloy must be mixed with 500 lb of a 22% aluminum alloy to make a 20% aluminum alloy?
Answer by ewatrrr(24785)


Hi
Let x represent the lbs of 15% aluminum
Question states***
'Mixing' to form (500+x)lbx of 20%
.15x + .22*500 = .20(500+x)
Solving for x
.02*500= .05x
.02*500/.05= x =200lb, the amount of 15% aluminum
CHECKING our Answer***
30 + 110 = .20*700 = 140